A horizontal pipe is 2 cm in radius at one end and increases in size so that it is 5 cm in radius at


the other end. Water is pumped into the small end of the pipe at 10 m/s and a pressure of


300 kPa (kilopascal). Find the


a.) speed and


b.) pressure of the water at the pipe’s large end.

1
Expert's answer
2021-12-06T09:40:50-0500

Explanations & Calculations


  • Use the continuity equation to calculate the speed at the other end.

"\qquad\qquad
\begin{aligned}
\small A_1v_1&=\small A_2v_2
\end{aligned}"

  • Calculate the are of the pipes at both ends in SI units: square meters.
  • Then apply them to the above equation, the speed at the other end can then be found.


  • Use Bernoulli's equation to calculate the pressure at the other end.

"\qquad\qquad
\begin{aligned}
\small P_1+\frac{1}{2}\rho v^2_1&=\small P_2+\frac{1}{2}\rho v_2^2
\end{aligned}"

  • Here "\small \rho" is the density of the fluid, water in this case.
  • You know by now the speed at both ends. Substitute the known values accordingly and get the answer for the pressure at the other end.